### v1

# Chain-of-Thought Prompting Obscures Hallucination Cues in Large Language Models: An Empirical Evaluation

Abstract
Large Language Models (LLMs) often exhibit hallucinations, generating factually incorrect or semantically irrelevant content in response to prompts. Chain-of-Thought (CoT) prompting can mitigate hallucinations by encouraging step-by-step reasoning, but its impact on hallucination detection remains underexplored. To bridge this gap, we conduct a systematic empirical evaluation. We begin with a pilot experiment, revealing that CoT reasoning significantly affects the LLM’s internal states and token probability distributions. Building on this, we evaluate the impact of various CoT prompting methods on mainstream hallucination detection methods across both instruction-tuned and reasoning-oriented LLMs. Specifically, we examine three key dimensions: changes in hallucination score distributions, variations in detection accuracy, and shifts in detection confidence. Our findings show that while CoT prompting helps reduce hallucination frequency, it also tends to obscure critical signals used for detection, impairing the effectiveness of various detection methods. Our study highlights an overlooked trade-off in the use of reasoning. Code is publicly available at: https://anonymous.4open.science/r/cot-hallu-detect.
Chain-of-Thought Prompting Obscures Hallucination Cues in Large Language Models: An Empirical Evaluation
Jiahao Cheng1  Tiancheng Su1  Jia Yuan1  Guoxiu He1 Jiawei Liu2Xinqi Tao3Jingwen Xie3Huaxia Li3 1 East China Normal University, Shanghai, China 2Wuhan University, Wuhan, China 3Xiaohsongshu Inc., Shanghai, China chengjiahao@stu.ecnu.edu.cn, gxhe@fem.ecnu.edu.cn
1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a range of natural language processing (NLP) tasks (Qin et al., 2024; Yan et al., 2024), showing a strong ability to follow instructions and answer questions (Zhao et al., 2024; Hadi et al., 2024). However, LLMs may also exhibit hallucinations. Specifically, they may produce factually incorrect or semantically irrelevant answers to a given prompt without indicating any lack of confidence (Ji et al., 2023; Yang et al., 2023; Hadi et al., 2024).
To reduce hallucinations in LLMs, extensive research has focused on hallucination detection and mitigation methods (Ji et al., 2023; Huang et al., 2024; Tonmoy et al., 2024). Existing detection methods are built upon various aspects of LLMs, including output consistency (e.g., evaluating the consistency of responses to multiple identical queries) (Wang et al., 2023; Manakul et al., 2023), analysis of internal states (e.g., computing confidence metrics) (Chuang et al., 2024; Chen et al., 2024d; Sriramanan et al., 2024), and self-evaluation of knowledge ( e.g., querying its certainty) (Kumar et al., 2024; Diao et al., 2024). Regarding hallucination mitigation, chain-of-thought (CoT) prompting improves response reliability and interpretability by guiding LLMs to generate intermediate reasoning steps before arriving at a final answer (Wei et al., 2022; Kojima et al., 2024). In contrast to more complex solutions such as fine-tuning LLMs or improving decoding strategies, CoT prompting is widely adopted due to its simplicity and effectiveness (Liu et al., 2023; Yang et al., 2024a). The performance improvements shown by GPT-4o OpenAI (2024) and DeepSeek-R1 Guo et al. (2025b) through enhanced reasoning have further fueled growing research interest in reasoning-centric LLMs.
However, as illustrated in Figure 1, the use of CoT prompting may diminish the effectiveness of hallucination detection methods. When an LLM is prompted to perform step-by-step reasoning, it semantically amplifies the LLM’s internal confidence in its output. As a result, even when the final answer deviates from the ground truth, the LLM tends to produce incorrect tokens with high confidence (i.e., high probability). Consequently, hallucinations induced by CoT prompting often appear more plausible, making them more difficult for detection methods to identify.
Figure 1: An example of CoT prompting confusing hallucination detection. When the input When is the Earth closest to the Sun? is provided to the LLM, it generates a hallucinatory response (during the summer) without CoT prompting. In this case, the hallucination detection method assigns a high hallucination score. However, after incorporating CoT prompting, where the LLM introduces reasoning steps (e.g., Step 1: Earth’s Orbit…), the LLM still produces the same incorrect response, but the detection method assigns only an average hallucination score.
Although some negative cases suggest that reasoning induced by CoT prompting may adversely affect hallucination detection, existing studies have yet to establish the generality of this phenomenon. To bridge this gap, we conduct a series of empirical evaluations. We begin with pilot experiments on three multiple-choice question answering (MCQA) tasks, including CommonSenseQA, ARC-Challenge, and MMLU, where we observe that CoT prompting significantly alters the probability distribution over the final answer tokens. Building on these findings, we perform a more extensive set of comparative experiments using two representative types of datasets: (1) complex fact-based question answering datasets, including TriviaQA, PopQA, HaluEval, and TruthfulQA; and (2) a summarization dataset requiring deep contextual understanding, named CNN/Daily Mail. We also incorporate four well-established, recently proposed hallucination detection methods to identify hallucinations. Furthermore, we evaluate three mainstream CoT prompting strategies across both instruction-tuned and reasoning-oriented LLMs, including LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, LLaMA-3.1-80B-Instruct, and DeepSeek-R1-Distill-Llama-8B. To comprehensively assess the impact of CoT prompting on hallucination detection, we examine three key dimensions: (1) shifts in the distribution of hallucination detection output scores; (2) changes in detection accuracy; and (3) variations in the confidence levels of the detection methods.
Our experimental results reveal a dual effect of CoT prompting: while it enhances LLM performance, it simultaneously disrupts hallucination detection mechanisms. This disruption arises because CoT prompting systematically alters the distribution of hallucination scores, leading to a counterintuitive trade-off. Particularly, improved task performance is accompanied by reduced hallucination detectability. The extent of this effect varies across detection paradigms: consistency-based methods exhibit greater robustness, whereas internal-state-based and self-evaluation-based methods are more susceptible to reasoning. Notably, advanced detection methods lose their typical advantage over perplexity-based baselines under CoT prompting. These findings suggest that reasoning-enhanced inference reshapes the landscape of hallucination detection, making hallucination signals more subtle and harder to capture.
Our contributions are threefold: (1) We conduct a series of experiments, including pilot and comprehensive evaluations, to empirically investigate the impact of CoT-induced reasoning on hallucination detection. (2) We reveal a double-edged effect of CoT prompting: while it enhances LLM performance, it simultaneously obscures common hallucination features. (3) Our results show that internal state-based and self-evaluation-based methods exhibit reduced robustness under CoT prompting.
2 Preliminary
CoT Prompting Methods
In this study, we focus on three representative methods: Zero-shot Chain of Thought (CoT, Kojima et al. (2024)), Least-to-Most (LtM, Zhou et al. (2023)), and Minimum Reasoning Path Prompting (MRPP, Chen et al. (2024c)). Detailed descriptions of these methods are provided in Appendix C.
Unless otherwise specified, all CoT prompting methods are implemented in a zero-shot configuration. This ensures that the LLM operates without access to task-specific examples, allowing us to isolate the effect of reasoning strategies. Following prior work, we adapt our objectives by tailoring prompt templates to suit each experimental setup.
Hallucination Detection Methods
We focus on hallucination detection methods that do not rely on external knowledge bases (e.g., fact-checking (Li et al., 2024)) or require additional fine-tuning and specialized training (e.g., ITI (Li et al., 2023b) and SAPLMA (Azaria and Mitchell, 2023)). This choice is motivated by the limited generalization ability of such methods across tasks and domains. Instead, we examine four main categories of hallucination detection methods: consistency-based methods (e.g., SelfCheckGPT-nli (Manakul et al., 2023)); internal-state-based methods (e.g., Perplexity, In-Context Sharpness (Chen et al., 2024d), and LLM-Check (Sriramanan et al., 2024)); self-evaluation-based methods (e.g., Verbalized Certainty (Kumar et al., 2024)); and hybrid methods (e.g., INSIDE (Chen et al., 2024a) and SelfCheckGPT-Prompt (Manakul et al., 2023)), which combine features from multiple categories. For more detailed descriptions of these methods, please refer to Appendix D.
3 Pilot Experiment and Observation
Recent research on multiple-choice question answering (MCQA) tasks (Kumar et al., 2024; Chen et al., 2024d; Quevedo et al., 2024) has revealed a strong correlation between the likelihood of a token being selected by an LLM and the correctness of the final answer. Building on these insights, we begin with a pilot experiment on token-level MCQA to examine how CoT-prompted reasoning influences the probability distribution over answer tokens. We pay particular attention to cases where incorrect answers or hallucinations persist despite the use of CoT prompting.
3.1 Pilot Experimental Setting
Datasets. We use three widely adopted MCQA datasets in our study. CommonsenseQA (validation split, 1,221 samples; Talmor et al. (2019)) is a benchmark designed to evaluate commonsense reasoning; we use its validation split since the test set lacks ground-truth answers. AI2 ARC (ARC-Challenge, test split, 1,172 samples; Boratko et al. (2018)) is a challenging dataset that focuses on science-related questions requiring advanced reasoning. MMLU (test split, 14,042 samples; Hendrycks et al. (2021)) is a large benchmark that spans a broad range of academic and professional domains, providing a comprehensive assessment of general knowledge and reasoning ability.
LLMs. We conduct the pilot experiments using three open-source LLMs, covering both instruction-tuned and reasoning-oriented variants: Llama-3.1-8B-Instruct (Dubey et al., 2024), a large-scale model fine-tuned for instruction following; Mistral-7B-v0.3-Instruct (Jiang et al., 2023), a compact model trained via instruction tuning; and DeepSeek-R1-Distill-Llama-8B Guo et al. (2025b), a reasoning-oriented model distilled from larger LLMs.
Metrics. We evaluate the LLMs using the following three metrics: Accuracy, which measures overall task performance by assessing the LLM’s ability to predict the correct option; Entropy, which quantifies the uncertainty in the LLM’s probability distribution over answer options, reflecting its confidence in decision-making; and AUROC (Area Under the Receiver Operating Curve), which treats answer correctness as a proxy hallucination label and evaluates how well the token-level probability distribution distinguishes between hallucinated and non-hallucinated responses.
The implementation details of the pilot experiment are available in Appendix F.
3.2 Observations
The experimental results are shown in Table 1. Overall, LLMs prompted with CoT consistently achieve higher accuracy across all datasets. For example, with Llama-3.1-8B-Instruct, CoT improves accuracy from 90.5, 90.16, and 80.79 to 97.67, 94.96, and 91.57 on the three datasets, respectively. However, the effectiveness of more advanced CoT variants is less consistent. While both LtM and MRPP employ multi-path reasoning, MRPP further incorporates confidence-based validation. As a result, LtM may suffer from ambiguity when faced with multiple plausible reasoning paths, achieving better performance than the base Llama but falling short of standard CoT. In contrast, MRPP outperforms even the basic CoT approach.
Table 1: Performance comparison of LLMs under CoT prompting on three token-level MCQA tasks. Experimental results show that CoT prompting improves accuracy and reduces entropy, suggesting increased confidence in the LLM’s predictions. However, AUROC scores also decline, indicating that token-level probabilities become less reliable as indicators of hallucinations. Detailed results for Mistral-7B-v0.3-Instruct are provided in the Appendix.
ARC-Challenge CommonSenseQA MMLU
Acc Entropy AUROC Acc Entropy AUROC Acc Entropy AUROC
DeepSeek-R1-Distill-Llama-8B
base 67.41 67.70 78.15 59.26 82.56 73.11 54.62 80.74 71.78
LtM 87.36 5.25 69.37 73.52 6.22 69.26 76.92 11.10 68.33
MRPP 88.05 6.93 72.90 72.62 7.13 67.44 76.54 12.08 67.52
CoT 87.94 4.86 70.15 72.38 5.45 68.34 77.36 10.56 64.23
Llama-3.1-8B-Instruct
base 90.50 23.96 85.98 90.16 26.33 79.75 80.79 46.94 79.97
LtM 93.42 18.14 78.07 89.93 27.82 76.17 85.67 36.31 78.62
MRPP 97.70 6.39 71.81 96.50 9.90 73.15 93.99 15.56 72.54
CoT 97.67 6.63 77.39 94.96 13.50 74.69 91.57 21.39 76.07
Figure 2: Comparison of token probability distributions for DeepSeek-R1-Distill-Llama-8B on ARC-Challenge. The upper row shows all samples; the lower row shows incorrect predictions after CoT prompting. Kolmogorov-Smirnov test confirms significant differences ($p<0.01$) between conditions with and without CoT prompting.
In addition, entropy consistently decreases across all LLMs after CoT prompting, suggesting that LLMs become more confident in their predictions, assigning higher probabilities to their selected answers. For instance, Llama’s entropy drops from 23.96, 26.33, and 46.94 to 6.63, 13.50, and 21.39, respectively, after applying CoT. Similar reductions are observed with other CoT strategies, reflecting a general improvement in model confidence.
The DeepSeek-R1-Distill-Llama-8B model, which is specifically optimized for reasoning abilities, performs substantially below expectations when reasoning is disabled. Its accuracy on the three datasets drops to 67.41, 59.26, and 54.62, respectively. When CoT is enabled, its performance improves significantly, though it still lags behind Llama in terms of accuracy. Interestingly, under CoT prompting, DeepSeek exhibits lower entropy than Llama across all datasets, suggesting greater confidence in its predictions, even though it produces more incorrect answers.
To better illustrate how confident LLMs are when producing hallucinated answers, i.e., incorrect predictions, we compare their token-level probability distributions before and after applying CoT prompting, for both all samples and hallucinated samples. The results are presented in Figure 2. Before CoT prompting, LLMs show relatively uniform probability distributions across samples; for hallucinated samples, the distributions are clearly left-skewed, indicating that incorrect answers tend to receive lower probabilities. In contrast, after CoT prompting, the distributions shift, and LLMs assign similarly high probabilities to both all and incorrect answers. This increased confidence, especially in wrong predictions, underscores the complex influence of CoT prompting on token-level probabilities, ultimately making hallucinations harder to detect.
Figure 3: The workflow for evaluating the impact of CoT prompting on hallucination detection. The upper section presents an overview, while the lower section highlights key implementation details. Given the LLM’s output, hallucinations are identified by comparing against ground-truth answers. A hallucination detection model then generates a score, which is used to produce predicted labels. Finally, detection performance is evaluated across four dimensions defined in the evaluation framework.
Previous studies have shown that the token probabilities produced by LLMs do not always reliably indicate hallucinations (Mahaut et al., 2024; Lin et al., 2022a; Chen et al., 2025). More robust signals can instead be extracted from the internal states of LLMs to better capture hallucination patterns. To explore this, we conduct a more comprehensive set of experiments using recent hallucination detection models, investigating whether CoT prompting influences the clarity of hallucination-related signals during the reasoning process.
4 Main Experiments
4.1 Pipeline
This section outlines the workflow used to assess whether CoT prompting interferes with hallucination detection, as illustrated in Figure 3. We begin by determining whether the LLM-generated response constitutes a hallucination based on the ground-truth response. Next, we compute a hallucination score using a detection model to predict whether the output is a hallucination. Finally, we examine the impact of CoT-generated reasoning on hallucination detection from three perspectives, including four evaluation metrics. Details are introduced below.
Correctness Measurement: For the responses generated by the LLM, we evaluate their alignment with the ground truth in the dataset based on text matching and semantic consistency. This evaluation determines whether a response constitutes a hallucination.
Hallucination Detection Score: During the response generation process, we apply four mainstream hallucination detection methods to compute the likelihood score of hallucination for each response. Based on these scores, we further predict whether a given response constitutes a hallucination.
Evaluation Dimensions:
Based on the hallucination labels and detection scores, we analyze the behavior of hallucination detection methods from surface-level patterns to deeper insights along three dimensions: score distribution, detection performance, and method confidence. Specifically: (1) changes in the distribution of hallucination scores are evaluated using a distribution consistency test; (2) detection performance is assessed via AUROC and the monotonic relationship; and (3) method confidence is measured using Expected Calibration Error (ECE). Further details are provided in Appendix E.
4.2 Experimental Setup
Datasets. We use four widely adopted QA datasets that require complex reasoning, along with one summarization dataset that demands strong in-context understanding. Specifically, we include TruthfulQA (Lin et al., 2022b), a dataset explicitly designed for hallucination detection, with a validation subset comprising 817 questions. In addition, TriviaQA (Joshi et al., 2017) is a large-scale reading comprehension dataset. We utilize its rc.wikipedia.nocontext and validation subsets. PopQA (Mallen et al., 2023) contains questions targeting a long-tail distribution of entities. We use its test subset in a closed-book setting. HaluEval (Li et al., 2023a) evaluates the factual accuracy of LLM-generated content by distinguishing between hallucinatory and accurate outputs. We use its QA subset of non-hallucinatory samples as a benchmark for correctness. For summarization, we employ the CNN/Daily Mail dataset Hermann et al. (2015), where the LLM is required to generate a summary based on a given news article. This task emphasizes the need for faithful summarization without deviation from the source content.
LLMs. Since some hallucination detection methods require access to the internal states of LLMs, we conduct our experiments using open-source LLMs. To ensure diversity across model families, parameter scales, and training objectives, we select four LLMs, including Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Llama-3.1-80B-Instruct, and DeepSeek-R1-Distill-Llama-8B.
The implementation details of the experiment can be found in Appendix F.
5 Results and analysis
5.1 Performance Comparison of CoT Methods Across Different Datasets
As shown in Table 2, consistent with the results of our pilot experiments, the CoT prompting methods significantly enhance the performance of LLMs across various tasks. By encouraging LLMs to explicitly generate intermediate reasoning steps, CoT prompting helps activate their latent knowledge and reasoning capabilities, leading to more accurate responses.
Notably, LLM performance rankings vary between the two evaluation approaches, textual matching and semantic consistency, highlighting the challenge posed by the diverse linguistic expressions in text generation. While evaluation methodology is not the primary focus of this study, to mitigate ambiguity, we adopt both metrics in subsequent experiments to assess response correctness, which serves as the ground truth for hallucination detection.
Furthermore, not all CoT variants consistently improve performance, suggesting that the reasoning process may have nuanced effects on output quality. Interestingly, DeepSeek performs better on summarization tasks, which demand stronger contextual understanding. This observation further supports the notion that incorporating reasoning can enhance LLM performance in such scenarios.
Table 2: Results of CoT prompting influencing the performance of LLMs across various datasets. The table presents results for Llama-3.1-8B-Instruct and DeepSeek-R1-Distill-Llama-8B on HaluEval, PopQA, TriviaQA, and CNN/Daily Mail (C./D.M.). The results indicate that CoT prompting substantially improves LLM performance on these tasks. Please refer to the appendix for additional tasks and LLMs.
HaluEval PopQA TriviaQA C./D. M.
Llama-3.1-8B-Instruct, Textual
base 30.26 30.31 79.41 22.13
CoT 36.59 35.05 81.60 23.33
LtM 36.98 33.57 80.02 22.72
MRPP 36.04 33.00 80.46 22.83
Llama-3.1-8B-Instruct, Semantic
base 39.89 48.58 85.61 91.80
CoT 77.37 65.30 85.17 91.07
LtM 38.86 46.79 85.63 90.79
MRPP 78.10 68.03 85.60 90.82
DeepSeek-R1-Distill-Llama-8B, Textual
base 19.19 16.09 61.78 20.04
CoT 29.21 22.09 63.16 22.54
LtM 30.15 19.01 64.10 21.39
MRPP 29.21 18.79 64.34 21.68
DeepSeek-R1-Distill-Llama-8B, Semantic
base 62.30 55.47 72.28 91.26
CoT 78.53 63.15 80.90 92.35
LtM 79.57 64.10 83.34 91.88
MRPP 79.35 64.34 83.02 91.36
5.2 Comparison of Score Distributions for Hallucination Detection
To examine the impact of CoT prompting on the internal decision-making processes of LLMs, we employ LLM-Check to compare the distributions of Hidden Score and Attention Score with and without CoT prompts. The Hidden Score reflects the LLM’s internal confidence in its own output, while the Attention Score captures the degree of focus allocated to different input segments.
We use the Kolmogorov–Smirnov (K–S) test to evaluate whether these distributions differ significantly across conditions. As shown in Figure 4, both scores undergo statistically significant shifts (p < 0.01), indicating that CoT prompts induce meaningful changes in the LLM’s internal representations.
Figure 4: Hidden score and attn score distributions for incorrect predictions. Each plot represents results across different CoT prompting methods (x-axis) and their corresponding scores (y-axis). Significant shifts in score distributions are observed for all hallucination detection methods after incorporating CoT prompting methods, suggesting that CoT prompting methods substantially impact hallucination detection scores: even when hallucinations exist, the scores of hallucination detection methods are biased towards not hallucination.
For instruction-tuned LLMs like LLaMA, the scores are relatively dispersed without CoT, suggesting that the hallucination detection methods can effectively differentiate between hallucinated and faithful responses. However, with CoT prompting, the scores become tightly clustered, implying that the detector may collapse to predicting a single class—thereby losing discriminative capability.
In contrast, reasoning-oriented LLMs such as DeepSeek exhibit concentrated score distributions even without CoT, suggesting that internal signals are less indicative of hallucination. CoT prompting still significantly alters the distribution, suggesting that explicit reasoning substantially reshapes the LLM’s internal state—even when the LLM is already strong in reasoning tasks.
Table 3: Number of experiments exhibiting AUROC decline across different LLMs and datasets.
HaluEval PopQA TriviaQA CNN/Daily Mail
Textual Semantic Textual Semantic Textual Semantic Textual Semantic
PPL 9 6 12 6 10 3 10 8
Sharpness 9 6 12 6 10 4 10 11
Eigen Score 3 8 4 7 9 6 4 7
SelfCk-Prompt 10 7 12 8 9 7 9 6
SelfCk-NLI 9 6 6 2 12 9 8 4
Hidden Score 8 9 4 6 8 5 6 7
Attn Score 5 4 0 7 1 5 12 7
Overall, these findings suggest that CoT prompting not only enhances output quality but also modulates the LLM’s internal confidence and attention.
5.3 Performance Comparison of Hallucination Detection Methods
We conduct a comprehensive evaluation of CoT prompting affects the performance of hallucination detection methods. Our experiments span three CoT prompting methods, four benchmark datasets, four LLMs, two hallucination annotation protocols (i.e., methods for establishing ground truth), and eight representative detection approaches, yielding a total of 768 experimental configurations. Table 3 summarizes the proportion of settings in which the AUROC score declined after applying CoT prompting. Overall, more than half of the setups (440 out of 768) exhibit performance degradation, suggesting that CoT prompting exerts a broad negative impact on hallucination detection. Notably, we treat TruthfulQA separately due to its distinct evaluation protocol.
Our findings reveal that CoT prompting generally impairs the effectiveness of hallucination detection, with the degree of impact varying by method. Self-evaluation-based approaches (e.g., SelfCheckGPT-Prompt and Verbalized Certainty) are most adversely affected. These methods rely heavily on confidence signals present in the LLM’s output, which are easily distorted by CoT-style reasoning. CoT prompting often causes LLMs to generate additional justificatory content that obscures factual inaccuracies. For instance, in response to the TriviaQA question What was Walter Matthau’s first movie?, an LLM under LtM prompting may include phrases like Based on reliable sources such as IMDb and Wikipedia…. Although the answer is incorrect, the Verbalized Certainty score may increase from 0.2 to 1.0, illustrating how CoT prompting can mask hallucinations.
By contrast, consistency-based methods (e.g., EigenScore and SelfCheckGPT-NLI) show greater robustness to CoT prompting. These methods leverage multi-sample reasoning to identify inconsistencies, which remain informative even in the presence of CoT prompts. In the above example, consistent sampling occasionally yields the correct answer The Kentuckian (EigenScore: 3/15; SelfCheckGPT-NLI: 1/20), albeit infrequently. While often criticized for high computational overhead, these approaches demonstrate strong resilience and reliability in reasoning-heavy scenarios.
A further notable observation is that although advanced hallucination detection methods typically outperform simple perplexity-based baselines under standard conditions, this advantage diminishes with CoT prompting. Without CoT, perplexity-based methods consistently underperform, aligning with prior studies (Chen et al., 2024a; Sriramanan et al., 2024). However, when CoT prompts are introduced, many detection methods fail to surpass the perplexity baseline, and some perform worse. This finding underscores a critical limitation of current detection approaches: they are not yet equipped to handle the structural and discourse-level shifts induced by explicit reasoning, ultimately reducing their ability to detect hallucinations effectively.
TruthfulQA. As shown in Table 4, CoT prompting exhibits inconsistent effects on enhancing the truthfulness and informativeness of responses in the TruthfulQA task across different scenarios (see Appendix G for details). Meanwhile, the negative impact of CoT prompting on hallucination detection methods varies significantly across different LLMs. Specifically, Llama shows a more pronounced performance drop in informativeness, Mistral experiences a greater decline in truthfulness, and DeepSeek shows a decline in both dimensions. This variation likely reflects differing response generation tendencies among LLMs. For example, after applying CoT prompting, Llama tends to emphasize the reasoning process and often includes redundant phrases such as Based on the above steps… in its answers. While such content highlights the reasoning chain, it may obscure factual inaccuracies or introduce misleading elements, thereby interfering with the judgment of hallucination detection methods.
Table 4: Number of experiments in which the AUROC value decreases. For each experiment, AUROC is calculated by comparing the predicted hallucination detection scores with either informativeness (I.) or truthfulness (T.); each cell shows the number of experiments (three using CoT prompting and one without) where AUROC decreases relative to a baseline. The first two columns are results for Llama-3.1-8B-Instruct, the middle two for Mistral-7B-Instruct-v0.3, and the last two for DeepSeek-R1-Distill-Llama-8B.
I. T. I. T. I. T.
Perplexity 3 0 0 2 2 3
Sharpness 3 0 0 2 2 3
Eigen Score 2 3 1 1 0 3
SelfCk-Prompt 1 3 1 0 2 3
SelfCk-NLI 1 2 0 3 2 3
Verb. Certainty 2 2 0 2 0 0
Hidden Score 2 0 3 3 2 2
Attn Score 0 2 3 3 1 3
Monotonic Relationship. The trends observed in monotonic relationships are consistent with those in classification performance. Comprehensive experimental results are provided in Appendix G.
5.4 Confidence in Hallucination Detection Methods
We further employ Expected Calibration Error (ECE) to evaluate the alignment between the hallucination probabilities predicted by the detection methods and the actual correctness of the hallucination detection, thereby assessing the calibration of model confidence. The results indicate that CoT prompting leads to an increase in ECE across most experimental settings, suggesting that the confidence of detection methods becomes less reliable. Detailed results are provided in Appendix G.
6 Discussion and Conclusion
In this study, we systematically investigate the impact of CoT prompting on hallucination detection in LLMs. We evaluate several representative detection methods and find that while CoT prompting improves text-level and semantic accuracy across most datasets, it simultaneously undermines detection performance. More specifically, CoT prompting leads to more concentrated detection score distributions, reduced classification accuracy, weaker alignment with ground truth, and poorer confidence calibration. The sensitivity of impact varies by method: self-evaluation-based approaches are most affected, whereas consistency-based methods exhibit greater robustness. We also observe that CoT prompting also diminishes, and occasionally eliminates, the performance advantage of advanced detectors over a simple perplexity baseline.
These findings suggest that CoT prompting can obscure hallucination cues and compromise existing detection systems, highlighting the need for more robust approaches tailored to reasoning-oriented LLM.
Limitations
Due to computational resource constraints, we are unable to conduct experiments on larger-scale LLMs, which limits the generalizability of our conclusions. Moreover, the limitations of current hallucination detection methods prevented us from evaluating our framework on widely used closed-source LLMs. Future research should incorporate such LLMs to achieve a more comprehensive understanding.
Regarding evaluation metrics, although the use of Exact Match and ROUGE-L to assess response correctness in QA tasks is subject to debate, we select these metrics because of their widespread adoption, computational efficiency, and effectiveness in measuring surface-level accuracy. Future work may consider more specialized metrics to better capture the nuanced challenges of hallucination detection.
Our findings show that while CoT prompting significantly improves the reasoning abilities of LLMs, it also introduces complexities that undermine existing hallucination detection mechanisms, posing risks in high-stakes applications. This underscores the urgent need to develop more robust hallucination detection methods tailored to reasoning-augmented LLM. Addressing the impact of CoT prompting on hallucination detection is beyond the scope of this study and is left for future work.
Ethics Statement
Our study investigates the impact of CoT reasoning on hallucination detection in LLMs. Importantly, this research does not involve the collection or use of personal data. However, we acknowledge the inherent risks associated with LLMs, including potential biases and misinformation. We are committed to conducting our research responsibly, ensuring transparency in our methodologies, and promoting the ethical use of AI technologies.
Appendix
Appendix A Theoretical and Practical Implications
The theoretical significance of this study lies in its exploration of the relationship between CoT prompting and hallucination detection. Our research demonstrates that while CoT prompting methods enhance the interpret-ability and reasoning capabilities of LLMs, they simultaneously confuse the task of detecting hallucinations in these LLMs. This highlights a gap in current research: hallucination detection methods must account for the effects introduced by CoT prompting. From a practical perspective, our work underscores the importance of considering a broader range of factors in hallucination detection. Specifically, we show that detection methods should be capable of handling the subtle differences introduced by CoT prompting. This is crucial for ensuring the safe and reliable deployment of LLMs in real-world applications, especially those requiring high levels of factual accuracy, such as healthcare, legal, and scientific fields.
Appendix B Additional Related Work
B.1 Large Language Models and Hallucinations
Large language models demonstrate remarkable generalization capabilities, leading to their widespread adoption across a variety of tasks (Zhao et al., 2024; Qin et al., 2024; Yan et al., 2024). They exhibit strong performance, especially in question-answering tasks (Zhao et al., 2024; Yang et al., 2023; Hadi et al., 2024). However, despite their effectiveness, LLMs are not infallible, and a significant challenge that persists is the phenomenon of hallucinations (Ji et al., 2023; Yang et al., 2023; Hadi et al., 2024).
Hallucinations in LLMs typically manifest when the model generates fabricated or inaccurate information in an attempt to answer a question (Yang et al., 2024b; Tonmoy et al., 2024). These hallucinations can be classified into three categories: Input-Conflicting, Context-Conflicting, and Fact-Conflicting hallucinations (Zhang et al., 2023). Alternatively, they are often categorized into two broad types: factuality hallucination and faithfulness hallucination (Huang et al., 2024). Factuality hallucinations arise when the LLM’s output conflicts with established world knowledge. Specifically, these hallucinations arise when the input prompt lacks relevant factual information. In such cases, the LLM relies on its internal parameters and memory (Mallen et al., 2023), resulting in outputs that contradict factual reality.
In the context of QA tasks, particularly closed-book QA, the issue of Fact-Conflicting or factuality hallucinations is especially pronounced (Jin et al., 2024; Zhang et al., 2023). Several benchmarks are designed to assess these types of hallucinations. Examples include TruthfulQA (Lin et al., 2022b) and HaluEval (Li et al., 2023a), which explicitly focus on hallucinations. Other QA datasets, such as TriviaQA (Joshi et al., 2017) and PopQA (Mallen et al., 2023), were not originally designed for hallucination detection but are still widely used to evaluate factual accuracy (Chen et al., 2024d, a, b).
A variety of methods are proposed to mitigate factuality hallucinations (Tonmoy et al., 2024), including prompt engineering, the integration of external knowledge, changes to decoding strategies, and model training or fine-tuning approaches. However, for users interacting directly with LLMs, prompt-based methods, such as Chain-of-Thought (Wei et al., 2022), provide a more accessible and widely adopted solution. The core idea behind CoT is to guide the LLM in generating reasoning steps before providing a final answer, thereby enhancing response accuracy and reducing hallucinations.
B.2 CoT prompting
Chain-of-Thought (CoT) prompting is a technique designed to enhance the performance of LLMs by guiding them to generate intermediate reasoning steps prior to producing a final answer (Wei et al., 2022). Rather than generating a direct one-step response, CoT prompts the LLM to engage in a structured reasoning process, thereby reducing errors and hallucinations that arise from premature conclusions. By leveraging this step-by-step reasoning approach, CoT demonstrates significant improvements in both accuracy and reliability.
To reduce the manual effort required to provide examples and to better leverage LLMs’ prior knowledge, Self-generated Few-shot CoT is introduced (Yasunaga et al., 2024). This approach allows LLMs to autonomously construct relevant examples and knowledge, enabling adaptive reasoning. Remarkably, studies demonstrate that performance improvements associated with reasoning can occur even without explicitly providing examples (Kojima et al., 2024). By simply appending a phrase such as “let’s think step by step” to the prompt — a technique known as Zero-shot CoT — LLMs are prompted to generate reasoning steps. This appended phrase (referred to as a “magic phrase” (Stechly et al., 2024)) is identified as a key trigger for enabling reasoning without explicit demonstrations.
Regardless of whether few-shot examples, self-generated examples, or zero-shot techniques are utilized, the primary objective of these methods is to prompt reasoning prior to delivering a final answer. This foundational idea predates the rise of LLMs (Nye et al., 2022; Ling et al., 2017). Beyond Zero-shot CoT, numerous variants emerge to address specific limitations. For example, Tree-of-Thought (ToT) (Yao et al., 2023) and Graph-of-Thought (GoT) (Yao et al., 2024) propose more sophisticated reasoning structures that facilitate backtracking and synergistic outcomes, thereby addressing the linear constraints of CoT reasoning. Other approaches, such as Least-to-Most Prompting (LtM) (Zhou et al., 2023) and Minimum Reasoning Path Prompting (Chen et al., 2024c), aim to decompose complex problems into simpler sub-problems or ensure that reasoning paths remain as concise and clear as possible.
Collectively, these methods share a common emphasis on generating reasoning steps prior to producing a final answer. For clarity, we refer to them as CoT prompting. They demonstrate broad utility, consistently enhancing LLM performance across various tasks and domains (Guo et al., 2025a; Kim et al., 2023; Nguyen et al., 2025). Substantial research examines both the empirical and theoretical effectiveness of CoT prompting (Feng et al., 2023; Chen et al., 2024c; Saparov and He, 2023; Prabhakar et al., 2024; Prystawski et al., 2023), establishing it as a robust strategy for performance improvement.
However, despite their widespread success, CoT prompting faces notable limitations (Sprague et al., 2024). Some studies argue that LLMs do not truly learn the underlying algorithms for solving problems (Stechly et al., 2024) and often fail to faithfully explain the reasoning behind their answers (Turpin et al., 2023). By requiring LLMs to generate additional content, such as reasoning steps, CoT prompting may alter the internal states of the LLM (Feng et al., 2023). This raises an intriguing question: could this state alteration impact tasks that rely on these internal states, such as hallucination detection?
B.3 Hallucination Detection
Hallucinations in LLMs are commonly attributed to a discrepancy between the LLM’s internal knowledge and real-world knowledge (Huang et al., 2024; Zhang et al., 2024). Some studies further suggest that hallucinations generated in earlier stages of text generation may propagate, leading to subsequent hallucinations (Zhang et al., 2024; Stechly et al., 2024). This observation aligns with our hypothesis that CoT prompting could influence hallucination detection by altering the LLM’s internal states.
Our findings focus on a different aspect: the impact of CoT prompting on hallucination detection. Specifically, even in cases where hallucinations do not result from the propagation of earlier errors, hallucination detection methods may exhibit different performance under CoT prompting compared to standard prompting strategies. This indicates that the additional reasoning steps introduced by CoT prompting may modify the internal states of model in ways that undermine the effectiveness of existing detection methods.
Methods for identifying and mitigating hallucinations can generally be classified based on whether they rely on external data. Among methods that do not rely on external data, three main subcategories are identified: consistency-based, internal-state-based, and self-evaluation-based hallucination detection.
Internal-state-based methods utilize the LLM’s internal representations or uncertainty during text generation to distinguish between hallucinated and non-hallucinated content. For instance, LLM-check (Sriramanan et al., 2024) assesses variations in token probabilities during text generation. Other approaches, such as DoLa (Chuang et al., 2024) and In-context Sharpness (Chen et al., 2024d), analyze features extracted from hidden states across different layers or time steps during decoding. These insights are then applied to detect and mitigate hallucinations.
Self-evaluation-based methods are founded on the premise that LLMs are capable of verbally evaluating the content they generate (Yao et al., 2023; Wang et al., 2023; Yao et al., 2024; Kumar et al., 2024; Diao et al., 2024). These methods employ prompts to extract the LLM’s self-evaluated confidence in its outputs. By treating the LLM’s verbalized certainty as an indicator of correctness, these methods identify hallucinations using confidence scores.
Consistency-based methods, such as SelfCheckGPT, operate under the assumption that if an LLM has accurate knowledge and confidence in its answers, it will generate semantically consistent outputs across multiple sampling attempts (Wang et al., 2023; Manakul et al., 2023). This approach is particularly relevant in factual QA tasks, where semantic consistency among sampled answers serves as an indicator of factuality. Inconsistent outputs, on the other hand, suggest the presence of hallucinations.
These three approaches are not mutually exclusive and can be combined. For example, INSIDE (Chen et al., 2024a) integrates internal-state-based and consistency-based methods, while SelfCheckGPT (Manakul et al., 2023) incorporates both self-evaluation-based and consistency-based techniques.
In contrast, hallucination detection methods that depend on external data are further categorized into retrieval-based and model-training-based approaches. Retrieval-based methods utilize external knowledge bases to validate generated outputs (Min et al., 2023; Sankararaman et al., 2024; Li et al., 2024). However, these methods often face limitations, including reliance on the quality and scope of external data, as well as the complexity of integrating retrieval and comparison systems. Model-training-based methods, such as SAPLMA (Azaria and Mitchell, 2023) and Inference-Time Intervention (ITI) (Li et al., 2023b), necessitate additional labeled data to train auxiliary models for hallucination detection. These approaches, while effective, are resource-intensive and often lack generalization ability. For instance, ITI not only identifies hallucinations but also employs trained probes to address them. Nevertheless, the high resource demands and restricted plug-and-play functionality render these methods less practical for direct application to existing LLMs.
In light of these challenges, our study concentrates on hallucination detection methods that do not rely on external data, with a particular emphasis on consistency-based, internal-state-based, and self-evaluation-based approaches. These methods provide enhanced flexibility and can be seamlessly integrated into existing LLM architectures without incurring additional resource overhead.
Appendix C Descriptions of CoT prompting
CoT encourages LLMs to think through problems step by step, fostering logical consistency and improving performance in complex tasks. In our experiments, we use the following adapted template: Think about it step by step.
LtM guides the LLM to decompose complex questions into simpler subproblems, solve them individually, and recombine the answers to address the original question. The template used in our study is: What subproblems must be solved before answering the inquiry?
MRPP aims to minimize cognitive load by ensuring that each reasoning step is as simple and error-free as possible. The adapted template we employ is: You need to perform multi-step reasoning, with each step carrying out as many basic operations as possible.
Given these challenges, our study focuses on hallucination detection methods that do not require external data, particularly consistency-based, internal-state-based, and self-evaluation-based approaches. These methods offer greater flexibility and can be readily applied to existing LLM architectures without additional resource overhead.
Appendix D Descriptions of Hallucination Detection Methods
Perplexity serves as a metric for assessing the predictive accuracy of an LLM on a given sequence. It is defined as:
$PPL(x)=2^{-\frac{1}{n}\sum_{i=1}^{n}\log P(x_{i}|x_{<i})}$ (1)
where $x=\{x_{1},x_{2},\dots,x_{n}\}$ denotes the token sequence, and $P(x_{i}|x_{<i})$ indicates the conditional probability of token $x_{i}$ given the preceding tokens.
In a well-trained language model, sequences that conform to its learned language patterns are assigned higher probabilities, thereby yielding lower perplexity. By contrast, hallucinated content frequently diverges from the statistical patterns present in the training data, resulting in lower token probabilities and higher perplexity. This renders perplexity a straightforward yet effective indicator for hallucination detection.
Sharpness (Chen et al., 2024d) represents an internal-state-based method designed to assess the concentration of a LLM’s hidden state during next-token prediction. The method introduces a metric termed contextual entropy, which measures the dispersion of the LLM’s internal activations across input context tokens. For a predicted token $v_{p}$, the contextual entropy is computed as:
$\begin{split}E(v_{p},v_{1:t})&=-\sum_{i=1}^{t}\tilde{P}(activation=i\mid v_{p},v_{1:t})\\ &\times\log\tilde{P}(activation=i\mid v_{p},v_{1:t}).\end{split}$ (2)
Here, $\tilde{P}(\text{activation}=i\mid v_{p},v_{1:t})$ denotes the normalized activation probability of token $v_{p}$ relative to each context token $v_{i}$. Lower entropy values signify more concentrated activations, implying sharper internal representations and greater confidence in token predictions.
In its original formulation, contextual entropy serves to adjust the original token probability distribution via a weighting coefficient. This adjustment enables the method to regulate the influence of entropy on token probabilities, thereby improving its capacity to detect factually incorrect predictions. By incorporating this entropy-based adjustment, the method successfully identifies discrepancies between high-confidence predictions and factually grounded content, rendering it well-suited for hallucination detection.
Verbalized Certainty (Kumar et al., 2024) is a self-evaluation-based method that enables the LLM to explicitly express its confidence level in the accuracy of its outputs. This is achieved by generating responses to a new query. The method employs prompts to elicit the LLM’s self-evaluation of response confidence, utilizing techniques such as Confidence Querying Prompts, Simulation of a Third-Person Perspective, and Likert Scale Utilization.
INSIDE (Chen et al., 2024a) represents a hybrid method that integrates internal-state-based and consistency-based approaches. It assesses self-consistency across multiple responses by calculating the EigenScore, which measures semantic diversity in the embedding space. For $K$ generated sequences, the covariance matrix of their sentence embeddings is computed as:
$\bm{\Sigma}=\mathbf{Z}^{\top}\cdot\mathbf{J}_{d}\cdot\mathbf{Z}$ (3)
where $\bm{\Sigma}\in\mathbb{R}^{K\times K}$ is the covariance matrix, $\mathbf{Z}=[\bm{z}_{1},\bm{z}_{2},\dots,\bm{z}_{K}]\in\mathbb{R}^{d\times K}$ represents the embedding matrix, and $\mathbf{J}_{d}=\bm{I}_{d}-\frac{1}{d}\mathbf{1}_{d}\mathbf{1}_{d}^{\top}$ is the centering matrix. The EigenScore is defined as the log-determinant of the regularized covariance matrix:
$E(\mathcal{Y}|\bm{x},\bm{\theta})=\frac{1}{K}\log\det(\bm{\Sigma}+\alpha\cdot\mathbf{I}_{K})$ (4)
where $\alpha\cdot\mathbf{I}_{K}$ ensures numerical stability, and the eigenvalues $\lambda_{i}$ are utilized to compute:
$E(\mathcal{Y}|\bm{x},\bm{\theta})=\frac{1}{K}\sum_{i=1}^{K}\log(\lambda_{i}).$ (5)
This metric effectively captures the self-consistency of generated responses and proves particularly adept at identifying hallucinations by detecting inconsistencies across multiple samples.
SelfCheckGPT (Manakul et al., 2023) utilizes consistency-based methods, including Natural Language Inference (NLI) and self-evaluation-based techniques, to evaluate hallucinations. In its NLI-based form, SelfCheckGPT compares LLM-generated responses with stochastically generated alternatives. It then categorizes the relationships between the hypothesis (response) and premise (context) into three categories: entailment, neutral, or contradiction. The contradiction score serves as the SelfCheckGPT score. Alternatively, in its self-evaluation-based form, SelfCheckGPT directly queries the LLM to verify whether specific sentences are supported by the generated context through carefully designed prompts.
LLM-Check (Sriramanan et al., 2024) represents an internal-state-based method that assesses hallucinations by analyzing the covariance structure of token embeddings within a single response. Similar to INSIDE’s EigenScore, LLM-Check computes the covariance matrix of hidden representations:
$\bm{\Sigma}_{2}=\mathbf{H}^{\top}\mathbf{H}$ (6)
and defines the hallucination score as the mean log-determinant of $\bm{\Sigma}_{2}$:
$\text{Score}=\frac{2}{m}\sum_{i=1}^{m}\log\sigma_{i}$ (7)
where $\sigma_{i}$ are the singular values of $\mathbf{H}$.
Appendix E Details on the Evaluation Framework
The evaluation framework aims to systematically examine the relationship between hallucination scores and correctness under both CoT prompting and non-CoT prompting conditions. The framework includes the following components:
Distribution Consistency Test: We use statistical methods, such as the Kolmogorov-Smirnov test, to compare the distribution of hallucination scores with and without CoT prompting. This analysis determines whether the introduction of CoT prompting significantly alters the distribution of hallucination scores. It also provides an intuitive measure of how CoT prompting influences hallucination detection.
Classification Performance: We treat correctness as the ground truth (binary: correct or incorrect) and hallucination scores as predictions. We then use AUROC to quantify the performance of hallucination detection methods before and after applying CoT prompting. This metric quantifies the degree to which CoT prompting influences the effectiveness of hallucination detection methods.
Monotonic Relationship: Using Spearman’s rank correlation coefficient, we evaluate the monotonic relationship between hallucination scores and correctness values. Here, correctness is treated as a continuous variable rather than a binary label. This metric shows whether hallucination detection scores are consistent with correctness under CoT prompting.
Alignment: We use Expected Calibration Error (ECE) to assess how well hallucination scores align with correctness. By treating correctness as the ground truth and hallucination scores as predictions, this metric evaluates whether hallucination detection methods effectively reflect the degree of correctness. This approach goes beyond simply distinguishing between correct and incorrect responses. This is particularly important as some answers may not be strictly correct or incorrect, and well-calibrated hallucination scores should reflect the nuances of correctness.
This evaluation framework collectively enables a comprehensive analysis of whether CoT prompting confuses hallucination detection methods. By observing performance degradation, changes in distribution, and alignment issues, our methodology provides a systematic approach to understanding and quantifying the impact of reasoning steps on hallucination detection.
Appendix F Experiments Implementation Details
F.1 Pilot Experiment
The normalization process for token probabilities in candidate options can be mathematically described as follows:
${p}_{\text{o}}=\frac{{p}_{\text{t}}[{t}_{\text{o}}]}{\sum_{i=1}^{N}{p}_{\text{t}}[{t}_{i}]},\hat{y}=\arg\max_{i}{p}_{\text{o}}[i],$ (8)
where ${t}_{\text{o}}$ and ${p}_{\text{o}}$ denote the tokens and normalized probabilities of each candidate option, respectively. $N$ is the total number of candidate options, and $\hat{y}$ is the LLM’s predicted option, determined by selecting the one with the highest probability. The predicted answer $\hat{y}$ is then compared to the ground-truth answer to determine whether the LLM’s response is correct. A correct prediction implies no hallucination, while an incorrect one corresponds to a potential hallucination.
For experiments involving CoT prompting, we first prompt the LLM to generate reasoning steps and then append a final query about the candidate options to determine the final answer. For the baseline (direct answer) setup, we directly prompt the LLM for the final answer without generating reasoning steps. We generate reasoning steps using the vLLM [^1] https://docs.vllm.ai/en/v0.6.1.post1/ with temperature = 0.5 and max_tokens = 512.
Once the complete prompt is constructed, we feed it into the LLM to compute token-level probabilities for the candidate options. These probabilities are recorded for subsequent calculations of accuracy, entropy, and AUROC.
By systematically comparing LLMs’ performance with and without CoT prompting, we aim to evaluate their ability to improve confidence and accuracy, as well as their impact on hallucination detection.
F.2 Main Experiment
[…]
We validate our automated metrics by manually annotating 150 samples with 2 annotators, achieving a Cohen’s Kappa of 0.863. Spearman’s $\rho$ between our automated scores and human annotations is 0.690 for Exact Match, 0.656 for ROUGE-L F1, and 0.683 for semantic similarity, confirming that our automated evaluations reliably reflect human judgment.
When generating reasoning content, we apply the corresponding prompts discussed in Section 2. For consistency-based hallucination detection methods such as INSIDE and SelfCheckGPT, which involve stochastic sampling, we adopt the generation parameters reported in their respective papers. For INSIDE, we set n=15, temperature=0.5, top_p=0.99, and top_k=5. For SelfCheckGPT, the parameters are n=20 and temperature=1.
For internal-state-based hallucination detection methods, including INSIDE, Sharpness, and LLM-Check, we determine layer-specific settings based on the best-performing configurations reported in the literature. Specifically, for INSIDE, we use layer 17; for Sharpness, layer 26; and for LLM-Check, we use layer 15 for the Hidden Score and layer 23 for the Attention Score. For Sharpness, which involves adjusting the entropy’s impact on the original token distribution using a parameter $\alpha$, we set $\alpha=1.0$ according to the settings in its original paper.
TruthfulQA offers a training set with 6.9k examples, which we use to fine-tune the GPT-3-6.7B model for evaluation purposes. Since OpenAI’s API no longer provides access to this model, we instead fine-tune the Llama-3.1-8B [^4] https://huggingface.co/meta-llama/Llama-3.1-8B model using LoRA (Hu et al., 2022) to compute the Truthfulness and Informativeness scores. We conduct fine-tuning on the LLaMA-Factory [^5] https://github.com/hiyouga/LLaMA-Factory platform with the following hyperparameters: learning rate = 5.0e-05, batch size = 32, epochs = 5, lora_rank = 8 and lora_alpha = 16.
The evaluation metrics are defined in Section 4.1. For Expected Calibration Error (ECE), we apply a normalization process to ensure predictions fall within the $[0,1]$ range. Specifically, we scale the data within the $3\%-97\%$ range to $[0,1]$, while outliers beyond this range are clipped to 00 or $1$. This normalization avoids distortions caused by extreme values.
Appendix G Results
Figure 5: Comparison of token probability distributions for Llama-3.1-8B-Instruct on ARC-Challenge. The upper row shows all samples; the lower row shows incorrect predictions after CoT prompting. Kolmogorov-Smirnov test confirms significant differences ($p<0.01$) between conditions with and without CoT prompting.
Table 5: Results of different LLMs and CoT prompting methods on the TruthfulQA dataset (values ×100). All metrics are positively oriented. Improvements after CoT prompting are bolded. Notably, the Reasoning model shows decreased Truthfulness but increased Informativeness after reasoning.
Truthfulness Informativeness
Llama-3.1-8B-Instruct
base 65.73 94.37
CoT 65.85 97.18
LtM 74.17 97.55
MRPP 52.63 95.23
Mistral-7B-Instruct-v0.3
base 85.68 93.64
CoT 87.88 97.06
LtM 80.29 95.72
MRPP 83.23 96.21
DeepSeek-R1-Distill-Llama-8B
base 56.18 94.86
CoT 21.18 97.92
LtM 25.83 96.45
MRPP 22.77 96.82
Table 6: AUROC results of Llama-3.1-8B-Instruct across different datasets. Bolded values indicate cases with reduced performance.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base 65.77 65.67 60.75 69.58 62.34 75.37 53.07 58.65
CoT 67.89 67.42 64.33 61.93 69.14 62.47 54.51 53.79
LtM 67.77 67.59 63.04 62.75 67.79 62.91 51.62 56.10
MRPP 67.48 67.22 66.73 62.82 67.75 61.54 57.34 56.46
Semantic HaluEval base 59.76 59.62 62.44 58.34 55.68 65.08 62.86 57.01
CoT 64.12 64.95 61.33 53.78 50.81 69.59 61.54 54.14
LtM 61.14 61.26 61.92 54.29 58.92 59.71 56.84 57.97
MRPP 70.33 69.63 64.20 51.10 62.77 77.51 55.96 61.79
Textual PopQA base 75.35 75.38 70.76 65.25 62.90 76.79 66.46 57.58
CoT 72.85 72.55 73.02 54.36 69.40 65.08 63.43 60.62
LtM 71.96 71.79 70.98 58.65 69.57 66.14 61.58 65.13
MRPP 68.73 68.53 69.91 60.17 68.56 62.29 63.52 61.29
Semantic PopQA base 71.90 72.13 72.16 53.05 51.68 80.03 72.07 66.24
CoT 66.06 66.11 66.49 54.77 56.23 64.90 61.96 57.28
LtM 73.60 73.62 72.58 56.70 71.80 67.08 64.15 67.50
MRPP 70.99 71.10 70.32 50.64 64.30 62.69 65.43 61.50
Textual TriviaQA base 75.56 75.30 70.39 77.20 82.21 64.39 59.43 55.13
CoT 69.86 69.53 67.62 69.32 77.87 59.39 57.24 56.86
LtM 70.46 70.34 68.07 66.64 71.61 58.60 58.77 61.49
MRPP 65.67 65.52 67.55 66.40 68.85 58.38 59.19 61.19
Semantic TriviaQA base 79.25 79.22 75.27 59.38 65.61 58.84 68.50 67.35
CoT 75.03 75.16 67.52 52.97 58.69 54.75 65.92 60.65
LtM 75.12 75.00 69.38 55.76 62.33 55.11 61.43 61.19
MRPP 79.46 79.18 74.27 63.69 63.86 56.95 67.12 60.78
Table 7: AUROC results of Mistral-7B-Instruct-v0.3 across different datasets. The bolded values indicate cases with reduced performance.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base 65.90 65.89 63.45 76.71 67.24 57.13 56.49 58.96
CoT 64.96 64.96 64.24 70.47 63.12 59.08 55.90 60.55
LtM 63.94 63.94 64.69 68.64 63.24 61.30 56.27 59.73
MRPP 64.18 64.18 64.96 69.11 59.26 62.46 59.47 61.39
Semantic HaluEval base 63.41 63.41 63.49 59.39 58.04 63.39 57.88 60.93
CoT 75.51 75.56 63.05 52.94 51.29 75.94 50.65 65.49
LtM 72.70 72.60 79.83 58.57 56.60 80.34 52.94 63.20
MRPP 76.65 76.74 73.10 51.89 63.33 76.63 51.81 69.32
Textual PopQA base 68.01 68.01 62.75 77.25 59.58 64.79 54.91 57.64
CoT 62.71 62.71 66.30 72.28 58.96 62.31 61.48 63.67
LtM 66.19 66.20 67.54 71.82 63.76 61.90 63.56 65.50
MRPP 60.99 61.00 68.64 72.11 55.24 67.01 65.51 67.00
Semantic PopQA base 59.55 59.56 56.97 65.23 60.73 59.62 58.84 59.87
CoT 58.47 58.47 54.32 54.29 60.26 58.30 53.42 51.63
LtM 63.73 63.73 63.43 60.68 64.21 55.92 61.86 63.93
MRPP 61.44 61.44 52.08 57.96 62.96 53.06 55.68 55.28
Textual TriviaQA base 64.17 64.20 60.37 84.50 73.47 59.00 55.29 57.75
CoT 66.57 66.81 67.77 76.36 68.47 57.95 60.53 62.07
LtM 62.56 62.65 67.28 74.20 68.51 58.40 62.90 63.45
MRPP 66.38 66.24 69.14 76.09 64.60 62.13 68.02 67.78
Semantic TriviaQA base 63.57 63.80 59.24 58.61 61.16 56.49 55.69 57.84
CoT 63.00 62.94 64.45 57.38 61.00 55.23 62.85 55.48
LtM 72.94 72.89 77.06 63.66 70.63 57.86 71.43 71.00
MRPP 71.62 71.61 71.45 65.30 67.96 61.09 65.55 61.25
Table 8: AUROC results of DeepSeek-R1-Distill-Llama-8B across different datasets. The bolded values indicate cases with reduced performance.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 0.5287 0.5281 0.6097 0.6464 0.6543 0.6094 0.6346
CoT 0.5142 0.5161 0.5192 0.6450 0.5932 0.5492 0.6156
LtM 0.5063 0.5089 0.5202 0.6276 0.5895 0.5117 0.6502
MRPP 0.5135 0.5129 0.5274 0.6313 0.5773 0.5338 0.6258
Semantic HaluEval base 0.6648 0.6643 0.6543 0.5033 0.5207 0.5382 0.7279
CoT 0.6546 0.6541 0.5513 0.5003 0.5493 0.5061 0.5058
LtM 0.6571 0.6563 0.5830 0.5491 0.5434 0.5019 0.5136
MRPP 0.6295 0.6323 0.5610 0.5389 0.5472 0.5073 0.5079
Textual PopQA base 0.5297 0.5288 0.5387 0.7073 0.6331 0.5134 0.5195
CoT 0.5096 0.5066 0.5372 0.6752 0.5874 0.5386 0.6479
LtM 0.5045 0.5051 0.5504 0.5869 0.5302 0.5647 0.5462
MRPP 0.5015 0.5055 0.5290 0.5659 0.5245 0.5507 0.5403
Semantic PopQA base 0.6584 0.6587 0.6052 0.5721 0.5124 0.5107 0.5606
CoT 0.6410 0.6380 0.6039 0.5019 0.5988 0.6202 0.5179
LtM 0.6010 0.5928 0.5547 0.5189 0.5768 0.5733 0.5319
MRPP 0.5770 0.5744 0.5457 0.5079 0.5514 0.5598 0.5300
Textual TriviaQA base 0.5417 0.5412 0.5716 0.7345 0.7132 0.5357 0.5315
CoT 0.5142 0.5125 0.5257 0.6255 0.5414 0.5362 0.6504
LtM 0.5179 0.5083 0.5178 0.6134 0.5565 0.5152 0.6684
MRPP 0.5244 0.5218 0.5150 0.6263 0.5431 0.5167 0.6592
Semantic TriviaQA base 0.5159 0.5155 0.5261 0.5602 0.6022 0.5857 0.5160
CoT 0.6588 0.6530 0.6027 0.5163 0.5337 0.5707 0.5352
LtM 0.6946 0.6915 0.6295 0.5175 0.5822 0.5937 0.5891
MRPP 0.7004 0.7002 0.6277 0.5286 0.5832 0.6171 0.6046
Textual CNN/Daily Mail base 0.6709 0.6697 0.5414 0.5077 0.5085 0.6315 0.6648
CoT 0.6133 0.6082 0.5757 0.5123 0.5035 0.5274 0.6043
LtM 0.6151 0.6079 0.5758 0.5002 0.5274 0.5576 0.6380
MRPP 0.6145 0.6082 0.5555 0.5068 0.5063 0.5428 0.6256
Semantic CNN/Daily Mail base 0.6769 0.6751 0.5557 0.5258 0.5210 0.5002 0.6716
CoT 0.5823 0.5734 0.5587 0.5075 0.5298 0.5143 0.6942
LtM 0.5919 0.5795 0.5610 0.5027 0.5457 0.5231 0.6877
MRPP 0.6004 0.5901 0.5798 0.5001 0.5242 0.5253 0.6998
Table 9: AUROC results of Llama-3.1-70B-Instruct across different datasets. Bolded values indicate cases with reduced performance.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base 0.7561 0.7561 0.7311 0.6877 0.6923 0.6182 0.7303 0.5721
CoT 0.6817 0.6802 0.7694 0.7014 0.6018 0.5503 0.6390 0.7005
LtM 0.7223 0.7226 0.7423 0.6503 0.6232 0.5542 0.6725 0.6847
MRPP 0.7491 0.7493 0.7762 0.6877 0.6250 0.5527 0.7451 0.6724
Semantic HaluEval base 0.7241 0.7234 0.6998 0.5782 0.5997 0.5938 0.6019 0.5165
CoT 0.6870 0.6870 0.6527 0.6226 0.5514 0.5275 0.6782 0.5889
LtM 0.7130 0.7128 0.7123 0.5816 0.5465 0.5664 0.6509 0.5980
MRPP 0.6805 0.6807 0.6586 0.5901 0.5653 0.5464 0.6269 0.5592
Textual PopQA base 0.7127 0.7126 0.6720 0.7293 0.6496 0.6773 0.6106 0.5185
CoT 0.6719 0.6717 0.7247 0.6923 0.6431 0.5762 0.6533 0.6393
LtM 0.6077 0.6078 0.6242 0.6052 0.5905 0.5426 0.5771 0.6089
MRPP 0.6908 0.6908 0.7261 0.6529 0.6721 0.5891 0.6437 0.6048
Semantic PopQA base 0.6497 0.6456 0.6578 0.5798 0.5916 0.5955 0.6694 0.5964
CoT 0.7363 0.7359 0.8020 0.7005 0.6449 0.5541 0.7513 0.6984
LtM 0.6295 0.6292 0.6467 0.5753 0.5743 0.5381 0.5932 0.6214
MRPP 0.7295 0.7289 0.7849 0.7035 0.6583 0.5553 0.7307 0.6519
Textual TriviaQA base 0.7225 0.7245 0.7337 0.6822 0.7610 0.7329 0.6212 0.6005
CoT 0.7041 0.7042 0.7039 0.7283 0.7090 0.6662 0.5961 0.5830
LtM 0.7024 0.7027 0.7085 0.7013 0.6804 0.6714 0.6183 0.6993
MRPP 0.7026 0.7026 0.6880 0.7325 0.7436 0.6836 0.6073 0.6591
Semantic TriviaQA base 0.7305 0.7305 0.7424 0.5611 0.6239 0.5866 0.5993 0.6542
CoT 0.7551 0.7551 0.7345 0.5200 0.5836 0.5055 0.7072 0.6766
LtM 0.7647 0.7650 0.7259 0.5902 0.6412 0.5582 0.5960 0.6465
MRPP 0.7394 0.7395 0.6958 0.5750 0.5927 0.5079 0.6574 0.5804
Table 10: AUROC results of CNN/Daily Mail dataset across different LLMs. Bolded values indicate cases with reduced performance.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Llama-3.1-8B-Instruct
Textual base 0.5891 0.5861 0.6295 0.5351 0.5877 0.5139 0.6332
CoT 0.5789 0.5791 0.6059 0.5001 0.5701 0.5386 0.6049
LtM 0.6267 0.6240 0.6444 0.5188 0.5376 0.5425 0.6138
MRPP 0.6049 0.5982 0.6228 0.5033 0.5517 0.5335 0.6040
Semantic
base 0.5681 0.5704 0.6395 0.5260 0.5610 0.5199 0.6695
CoT 0.5531 0.5510 0.5957 0.5278 0.5258 0.5144 0.6625
LtM 0.5873 0.5777 0.6421 0.5521 0.5172 0.5374 0.6677
MRPP 0.5559 0.5527 0.5858 0.5218 0.5089 0.5576 0.6499
Mistral-7B-Instruct-v0.3
Textual base 0.6219 0.6219 0.5960 0.5189 0.5314 0.6053 0.6404
CoT 0.5556 0.5556 0.5264 0.5081 0.5150 0.5404 0.6033
LtM 0.5785 0.5785 0.5583 0.5221 0.5294 0.5392 0.6061
MRPP 0.5564 0.5565 0.5216 0.5131 0.5303 0.5244 0.6043
Semantic base 0.6132 0.6132 0.5899 0.5256 0.5254 0.5249 0.7037
CoT 0.5560 0.5560 0.5296 0.5240 0.5708 0.5058 0.7155
LtM 0.5559 0.5559 0.5527 0.5470 0.5458 0.5141 0.6780
MRPP 0.5315 0.5315 0.5409 0.5202 0.5403 0.5243 0.7055
Llama-3.1-70B-Instruct
Textual base 0.5554 0.5569 0.5120 0.5612 0.5396 0.5157 0.5569
CoT 0.5203 0.5203 0.5446 0.5537 0.5473 0.5323 0.5221
LtM 0.5076 0.5074 0.5142 0.5451 0.5045 0.5978 0.5221
MRPP 0.5107 0.5111 0.5771 0.5891 0.5590 0.5769 0.5120
Semantic base 0.5588 0.5583 0.5455 0.5125 0.5261 0.6004 0.7052
CoT 0.5293 0.5289 0.5072 0.5177 0.5298 0.5893 0.6716
LtM 0.5216 0.5219 0.5062 0.5056 0.5248 0.5674 0.6806
MRPP 0.5558 0.5559 0.5563 0.5642 0.5649 0.5368 0.6449
Table 11: Spearman correlation results of Llama-3.1-8B-Instruct across different datasets. The bolded values indicate cases with reduced correlation.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual base -25.10 -24.94 -17.11 31.15 -19.64 46.46 4.89 13.77
CoT -29.85 -29.07 -23.91 19.91 -31.93 25.68 7.52 6.32
LtM -29.72 -29.42 -21.81 21.33 -29.75 26.53 2.71 10.20
MRPP -29.07 -28.64 -27.82 21.32 -29.53 24.15 12.21 10.74
Semantic base -19.59 -19.27 -28.14 14.93 -13.04 30.52 28.48 18.60
CoT -41.67 -41.99 -29.48 7.53 -10.86 29.39 27.54 6.48
LtM -21.30 -21.53 -23.95 8.13 -17.43 21.99 13.93 15.66
MRPP -45.41 -45.58 -35.17 17.97 -17.22 33.96 27.44 10.26
Textual base -40.36 -40.41 -33.04 24.28 -20.54 49.19 26.20 12.07
CoT -37.77 -37.27 -38.04 7.20 -32.07 28.92 22.20 17.55
LtM -35.93 -35.64 -34.32 14.15 -32.01 31.51 18.95 24.74
MRPP -30.51 -30.18 -32.44 16.57 -30.23 24.51 22.02 18.39
Semantic base -45.00 -45.28 -45.40 4.89 -7.46 62.95 41.78 37.74
CoT -47.24 -47.26 -47.26 5.45 -24.11 40.39 36.40 24.69
LtM -44.73 -44.92 -43.14 12.11 -41.46 39.02 26.11 33.81
MRPP -54.69 -54.79 -56.11 17.77 -43.34 37.45 41.24 32.82
Textual base -35.80 -35.44 -28.56 38.10 -45.12 40.24 13.21 7.18
CoT -26.66 -26.22 -23.66 25.94 -37.41 24.10 9.71 9.20
LtM -28.34 -28.18 -25.03 23.05 -29.93 24.13 12.14 15.92
MRPP -21.52 -21.31 -24.11 22.52 -25.89 23.30 12.62 15.38
Semantic base -57.87 -57.87 -47.63 17.37 -28.49 32.81 31.09 27.16
CoT -51.38 -51.71 -35.49 5.65 -17.42 19.28 30.22 19.60
LtM -53.37 -53.21 -40.77 13.00 -26.63 21.69 23.98 23.72
MRPP -59.79 -59.28 -48.05 29.86 -26.30 27.13 31.42 19.20
Table 12: Spearman correlation results of Mistral-7B-Instruct-v0.3 across different datasets. The bolded values indicate cases with reduced correlation.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base -27.09 -27.08 -22.92 45.51 -29.37 16.40 11.06 15.26
CoT -24.96 -24.97 -23.76 34.16 -21.90 17.88 9.84 17.62
LtM -23.03 -23.04 -24.26 30.79 -21.87 19.41 10.35 16.08
MRPP -23.08 -23.08 -24.34 31.10 -15.07 21.79 15.42 18.54
Semantic HaluEval base -21.87 -21.86 -22.16 16.47 -16.01 26.81 15.33 20.13
CoT -37.38 -37.38 -30.76 16.77 -9.00 38.16 16.26 1.65
LtM -55.59 -55.59 -53.57 34.42 -24.33 45.13 28.04 18.37
MRPP -38.08 -38.08 -36.07 24.07 -11.02 47.15 13.79 1.55
Textual PopQA base -29.69 -29.70 -21.02 44.93 -15.79 29.29 8.10 12.60
CoT -20.61 -20.62 -26.45 36.14 -14.54 23.23 18.63 22.17
LtM -26.11 -26.11 -28.28 35.18 -22.19 23.39 21.86 24.98
MRPP -17.05 -17.05 -28.91 34.28 -8.12 28.59 24.05 26.37
Semantic PopQA base -16.99 -16.99 -13.97 29.24 -22.63 15.92 13.69 19.34
CoT -28.93 -28.93 -30.22 28.55 -23.69 16.20 16.98 17.15
LtM -51.36 -51.37 -53.90 39.29 -41.04 27.23 40.45 47.09
MRPP -36.38 -36.39 -40.55 34.64 -27.78 32.72 20.07 20.00
Textual TriviaQA base -19.28 -19.33 -14.11 46.94 -31.94 20.66 7.20 10.55
CoT -23.90 -24.24 -25.62 38.01 -26.62 20.61 15.18 17.40
LtM -18.26 -18.40 -25.12 35.18 -26.91 23.42 18.75 19.55
MRPP -24.87 -24.66 -29.05 39.61 -22.17 26.29 27.36 26.99
Semantic TriviaQA base -27.44 -27.77 -17.31 13.55 -19.10 17.32 11.50 17.15
CoT -34.17 -34.25 -36.79 19.73 -29.70 18.67 34.88 16.45
LtM -50.34 -50.82 -56.95 30.10 -43.43 24.65 44.21 43.06
MRPP -51.60 -51.69 -49.91 34.31 -42.28 30.66 40.78 29.99
Table 13: Spearman correlation results of DeepSeek-R1-Distill-Llama-8B across different datasets. The bolded values indicate cases with reduced correlation.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score
Textual HaluEval base -0.0378 -0.0369 -0.1444 0.1928 -0.2031 0.1440 0.1772
CoT -0.0224 -0.0253 -0.0302 0.2285 -0.1468 0.0775 0.1821
LtM -0.0099 -0.0139 -0.0315 0.1994 -0.1398 0.0183 0.2347
MRPP -0.0212 -0.0203 -0.0430 0.2062 -0.1213 0.0530 0.1975
Semantic HaluEval base -0.3112 -0.3104 -0.2987 0.0689 -0.0066 0.0434 0.3913
CoT -0.3107 -0.3040 -0.1378 0.0165 -0.0120 0.0884 0.0135
LtM -0.3373 -0.3267 -0.1399 0.0554 -0.0557 0.0447 0.0510
MRPP -0.2884 -0.2892 -0.1580 0.0308 -0.0011 0.0686 0.0354
Textual PopQA base -0.0377 -0.0366 -0.0492 0.2638 -0.1694 0.0171 0.0249
CoT -0.0137 -0.0094 -0.0535 0.2518 -0.1256 0.0555 0.2126
LtM -0.0062 -0.0069 -0.0685 0.1181 -0.0411 0.0880 0.0628
MRPP -0.0020 -0.0075 -0.0393 0.0892 -0.0331 0.0686 0.0545
Semantic PopQA base -0.2394 -0.2408 -0.1381 0.1137 -0.0389 0.1260 0.0708
CoT -0.3860 -0.3809 -0.3123 0.0139 -0.2666 0.3350 0.0633
LtM -0.2425 -0.2308 -0.1677 0.0060 -0.1553 0.1619 0.0818
MRPP -0.2014 -0.1932 -0.1238 0.0183 -0.1239 0.1484 0.1111
Textual TriviaQA base -0.0722 -0.0714 -0.1241 0.4061 -0.3693 0.0619 0.0546
CoT -0.0238 -0.0209 -0.0430 0.2098 -0.0692 0.0605 0.2513
LtM -0.0301 -0.0139 -0.0300 0.1909 -0.0952 0.0255 0.2834
MRPP -0.0410 -0.0365 -0.0251 0.2121 -0.0724 0.0281 0.2672
Semantic TriviaQA base -0.0237 -0.0231 -0.0419 0.0866 -0.2094 0.2345 0.0126
CoT -0.3775 -0.3700 -0.2477 0.0498 -0.1374 0.1874 0.0859
LtM -0.4540 -0.4509 -0.3530 0.0348 -0.2267 0.2769 0.1660
MRPP -0.4320 -0.4251 -0.3374 0.0134 -0.2362 0.2881 0.1743
Textual CNN/Daily Mail base -0.3460 -0.3437 -0.0701 0.0140 -0.0183 0.2672 0.3268
CoT -0.2458 -0.2359 -0.1568 0.0055 -0.0327 0.0781 0.2400
LtM -0.2476 -0.2323 -0.1624 0.0011 -0.0478 0.1002 0.2848
MRPP -0.2552 -0.2445 -0.1238 0.0119 -0.0304 0.1083 0.2755
Semantic CNN/Daily Mail base -0.3661 -0.3631 -0.1145 0.0553 -0.0312 0.0001 0.3559
CoT -0.1898 -0.1693 -0.1297 0.0057 -0.0643 0.0427 0.3995
LtM -0.1963 -0.1744 -0.1141 0.0135 -0.1088 0.0483 0.3868
MRPP -0.1982 -0.1768 -0.1362 0.0121 -0.0649 0.0428 0.4083
Table 14: Spearman correlation results of Llama-3.1-70B-Instruct across different datasets. The bolded values indicate cases with reduced correlation.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base -0.4222 -0.4222 -0.4078 0.3499 -0.3632 0.2789 0.2739 0.0287
CoT -0.3453 -0.3435 -0.4282 0.3363 -0.1636 0.1409 0.2894 0.2698
LtM -0.3669 -0.3673 -0.3753 0.2490 -0.1571 0.1589 0.2312 0.2532
MRPP -0.3914 -0.3915 -0.4015 0.2812 -0.2027 0.1650 0.1693 0.2208
Semantic HaluEval base -0.4651 -0.4638 -0.4184 0.1525 -0.2041 0.2288 0.1875 0.1358
CoT -0.4201 -0.4201 -0.3420 0.2450 -0.1057 0.0682 0.3507 0.1896
LtM -0.4958 -0.4955 -0.4731 0.1732 -0.1472 0.1527 0.2608 0.2119
MRPP -0.4128 -0.4130 -0.3585 0.1997 -0.1415 0.1019 0.2616 0.1297
Textual PopQA base -0.3591 -0.3591 -0.2905 0.2583 -0.2526 0.3334 0.1868 0.3130
CoT -0.2914 -0.2912 -0.3811 0.3261 -0.2426 0.1416 0.2599 0.2361
LtM -0.1838 -0.1840 -0.2121 0.1796 -0.1544 0.0831 0.1316 0.1859
MRPP -0.3251 -0.3250 -0.3852 0.3906 -0.2931 0.1672 0.2448 0.1786
Semantic PopQA base -0.3100 -0.3014 -0.3241 0.1378 -0.2146 0.1757 0.3143 0.2033
CoT -0.4383 -0.4375 -0.5465 0.3819 -0.2936 0.1232 0.4460 0.3501
LtM -0.2768 -0.2764 -0.2985 0.1543 -0.1794 0.0887 0.1732 0.2345
MRPP -0.4190 -0.4180 -0.5272 0.3701 -0.2960 0.0957 0.4043 0.2700
Textual TriviaQA base -0.2103 -0.2121 -0.2208 0.1721 -0.2467 0.2355 0.1145 0.0950
CoT -0.1877 -0.1879 -0.1876 0.2100 -0.1922 0.1698 0.0884 0.0763
LtM -0.1930 -0.1933 -0.1989 0.1920 -0.1721 0.1910 0.1129 0.1901
MRPP -0.1882 -0.1882 -0.1746 0.2159 -0.2262 0.1983 0.0997 0.1478
Semantic TriviaQA base -0.3406 -0.3407 -0.4543 0.0085 -0.1633 0.1357 0.1771 0.2031
CoT -0.5110 -0.5108 -0.4573 0.1115 -0.1818 0.0083 0.3921 0.3450
LtM -0.5241 -0.5245 -0.3985 0.1840 -0.3022 0.1412 0.1937 0.2858
MRPP -0.4831 -0.4831 -0.3878 0.1527 -0.2058 0.0456 0.3112 0.1851
Table 15: Spearman correlation results of CNN/Daily Mail dataset across different LLMs. The bolded values indicate cases with reduced correlation.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Llama-3.1-8B-Instruct
Textual base -0.1595 -0.1591 -0.2852 -0.0788 -0.1843 -0.0274 0.2857
CoT -0.2116 -0.2067 -0.2452 -0.0039 -0.1306 -0.0777 0.2366
LtM -0.2725 -0.2664 -0.2993 -0.0438 -0.0853 -0.0885 0.2423
MRPP -0.2407 -0.2307 -0.2875 -0.0039 -0.1113 -0.0960 0.2392
Semantic
base -0.1582 -0.1620 -0.2949 -0.0687 -0.1287 -0.0897 0.3541
CoT -0.1192 -0.1136 -0.2083 -0.0488 -0.0448 -0.0454 0.3494
LtM -0.1941 -0.1761 -0.2796 -0.0880 -0.0339 -0.0348 0.3452
MRPP -0.1242 -0.1160 -0.1790 -0.0276 -0.0175 -0.1009 0.3150
Mistral-7B-Instruct-v0.3
Textual base -0.2385 -0.2385 -0.2061 0.0514 -0.0699 -0.2149 0.2910
CoT -0.1323 -0.1323 -0.0490 0.0032 -0.0384 -0.0641 0.2254
LtM -0.1821 -0.1822 -0.1269 0.0678 -0.0483 -0.0633 0.2154
MRPP -0.1140 -0.1141 -0.0376 0.0370 -0.0591 -0.0350 0.2280
Semantic
base -0.2227 -0.2226 -0.1985 0.0545 -0.0499 -0.0371 0.4040
CoT -0.1071 -0.1070 -0.0424 0.0389 -0.1306 -0.0056 0.4221
LtM -0.$1$0.1112 -0.1075 0.0860 -0.0802 -0.0061 0.3723
MRPP -0.0652 -0.0652 -0.0797 0.0270 -0.0762 -0.0292 0.4113
Llama-3.1-70B-Instruct
Textual base -0.1201 -0.1227 -0.0165 0.0837 -0.0609 -0.0335 0.1102
CoT -0.0026 -0.0017 -0.0785 0.1075 -0.0067 -0.0504 0.0221
LtM -0.0127 -0.0133 -0.0040 0.0686 -0.0481 -0.1914 0.0905
MRPP -0.0414 -0.0419 -0.1780 0.0521 -0.0263 -0.1630 0.0712
Semantic base -0.1044 -0.1033 -0.0924 0.0153 -0.0030 -0.2155 0.4319
CoT -0.0617 -0.0606 -0.0324 0.0429 -0.0079 -0.1417 0.3805
LtM -0.0053 -0.0053 -0.0269 0.0169 -0.0719 -0.1107 0.3714
MRPP -0.0732 -0.0729 -0.1472 0.0887 -0.0621 -0.0834 0.2947
Table 16: ECE results of Llama-3.1-8B-Instruct across different datasets. The bolded values indicate cases with reduced calibration.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 16.89 17.00 24.90 8.40 23.64 25.90 40.09
CoT 14.98 15.03 20.25 12.05 17.37 28.74 21.90
LtM 14.93 15.29 27.23 11.27 14.35 27.12 21.45
MRPP 14.95 15.31 24.70 11.10 15.64 28.64 18.63
Semantic HaluEval base 19.50 19.34 13.37 16.08 21.90 25.24 37.27
CoT 66.88 66.74 24.01 58.64 55.80 33.10 45.99
LtM 16.50 15.59 20.38 17.08 19.77 21.68 17.76
MRPP 65.31 65.12 20.52 63.78 64.02 34.32 50.81
Textual PopQA base 20.81 21.18 12.67 9.69 22.75 40.21 53.10
CoT 18.09 18.28 13.90 18.11 19.28 31.91 20.50
LtM 18.80 18.84 22.04 15.69 17.46 30.82 16.65
MRPP 18.86 18.82 29.71 12.37 18.40 33.39 19.04
Semantic PopQA base 15.29 15.61 4.98 16.66 27.78 27.39 41.15
CoT 46.31 45.67 21.58 44.41 43.24 21.19 34.10
LtM 15.25 15.06 14.73 15.83 18.23 24.49 11.62
MRPP 47.60 47.57 22.24 56.90 52.17 24.34 40.55
Textual TriviaQA base 23.98 24.30 27.40 40.62 22.30 10.65 10.93
CoT 27.32 27.81 23.50 42.41 23.45 16.46 29.36
LtM 29.77 29.29 26.69 40.20 27.11 17.08 25.70
MRPP 33.83 35.21 23.63 46.93 32.66 13.95 28.99
Semantic TriviaQA base 11.20 12.07 11.90 24.04 21.31 10.47 20.79
CoT 9.45 11.74 11.56 21.78 24.11 16.33 13.94
LtM 8.64 9.38 11.87 19.29 23.85 16.30 14.38
MRPP 12.08 12.23 13.44 23.48 20.60 13.34 15.41
Table 17: ECE results of Mistral-7B-Instruct-v0.3 across different datasets. The bolded values indicate cases with reduced calibration.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 12.80 12.75 23.43 19.43 17.78 18.71 23.51
CoT 17.06 17.10 26.82 38.75 19.96 18.93 14.34
LtM 18.26 18.24 29.60 40.78 18.42 20.62 14.58
MRPP 18.63 18.67 30.54 43.73 20.05 19.71 13.19
Semantic HaluEval base 11.26 11.23 19.60 26.37 23.74 14.24 18.15
CoT 56.92 56.91 17.20 22.25 61.50 47.64 56.51
LtM 60.89 60.89 25.75 21.92 67.90 46.80 56.03
MRPP 55.47 55.43 17.86 20.97 68.08 48.73 58.77
Textual PopQA base 17.76 17.74 26.49 22.52 28.24 21.89 25.30
CoT 19.12 19.15 22.59 40.93 20.86 19.84 14.76
LtM 21.59 21.61 31.86 42.11 18.28 18.43 11.04
MRPP 22.19 22.18 30.76 47.66 22.37 20.88 12.72
Semantic PopQA base 14.19 14.15 18.56 22.49 25.23 15.83 18.74
CoT 55.14 55.15 17.28 23.01 60.14 44.88 53.89
LtM 56.82 56.83 26.16 22.47 62.59 47.89 60.39
MRPP 51.42 51.44 15.97 21.37 65.40 49.18 60.06
Textual TriviaQA base 31.75 31.82 13.52 6.04 30.41 25.51 22.26
CoT 29.45 29.37 19.38 12.32 32.03 30.01 32.30
LtM 33.90 33.42 28.24 13.69 31.64 26.12 33.15
MRPP 26.24 26.14 23.55 15.48 33.93 21.80 31.19
Semantic TriviaQA base 11.45 11.98 23.09 36.09 25.21 14.18 14.59
CoT 21.23 21.51 15.00 27.66 35.64 18.23 24.38
LtM 26.65 26.63 25.76 16.83 40.52 25.51 33.13
MRPP 23.87 24.06 21.50 19.57 37.18 21.20 31.04
Table 18: ECE results of DeepSeek-R1-Distill-Llama-8B across different datasets. The bolded values indicate cases with reduced calibration.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score
Textual HaluEval base 0.2133 0.2098 0.4199 0.5159 0.3615 0.3567 0.4295
CoT 0.2088 0.2060 0.4256 0.6087 0.1700 0.3016 0.4706
LtM 0.1774 0.1789 0.4124 0.6148 0.1575 0.3186 0.4798
MRPP 0.2133 0.2053 0.4124 0.6052 0.1911 0.3225 0.4860
Semantic HaluEval base 0.2590 0.2607 0.1069 0.1943 0.2234 0.1299 0.1212
CoT 0.4416 0.4446 0.2364 0.0880 0.7213 0.2962 0.1452
LtM 0.4757 0.4789 0.2534 0.0805 0.7346 0.3158 0.1396
MRPP 0.4711 0.4763 0.2551 0.0817 0.7330 0.2981 0.1345
Textual PopQA base 0.2064 0.2054 0.4866 0.3780 0.4322 0.3659 0.5111
CoT 0.2726 0.2647 0.4598 0.6850 0.1289 0.3245 0.4400
LtM 0.3204 0.3138 0.5445 0.7279 0.1579 0.3846 0.4601
MRPP 0.3286 0.3196 0.5573 0.7268 0.1879 0.3744 0.4461
Semantic PopQA base 0.2273 0.2270 0.2293 0.2563 0.3350 0.2059 0.1713
CoT 0.1942 0.2071 0.1469 0.2231 0.4832 0.1389 0.1533
LtM 0.2247 0.2355 0.1348 0.2221 0.5359 0.1425 0.1501
MRPP 0.2445 0.2472 0.1421 0.2138 0.5690 0.1619 0.1721
Textual TriviaQA base 0.1205 0.1216 0.1923 0.2668 0.1416 0.1565 0.2142
CoT 0.2509 0.2548 0.2249 0.2752 0.4003 0.1632 0.1546
LtM 0.2644 0.2795 0.2367 0.2925 0.3937 0.1656 0.1626
MRPP 0.2697 0.2745 0.2267 0.2866 0.4178 0.1607 0.1509
Semantic TriviaQA base 0.2113 0.2074 0.2000 0.2514 0.2351 0.1268 0.1734
CoT 0.3014 0.3050 0.1707 0.1635 0.6069 0.2051 0.1071
LtM 0.3645 0.3699 0.2390 0.1186 0.6572 0.2456 0.0904
MRPP 0.3463 0.3470 0.2176 0.1219 0.6215 0.2232 0.0784
Textual CNN/Daily Mail base 0.1048 0.1093 0.3753 0.4203 0.3140 0.1259 0.2378
CoT 0.1534 0.1503 0.2742 0.2709 0.4309 0.1850 0.1279
LtM 0.1149 0.1144 0.2721 0.3257 0.3285 0.1443 0.1390
MRPP 0.1232 0.1176 0.3332 0.3311 0.3695 0.1538 0.1513
Semantic CNN/Daily Mail base 0.0950 0.0969 0.2595 0.3013 0.4183 0.1774 0.1289
CoT 0.1321 0.1298 0.4058 0.3618 0.3593 0.1851 0.1649
LtM 0.1184 0.1227 0.4462 0.3650 0.3173 0.1869 0.1987
MRPP 0.1179 0.1286 0.3661 0.4099 0.3243 0.1781 0.1988
Table 19: ECE results of Llama-3.1-70B-Instruct across different datasets. The bolded values indicate cases with reduced calibration.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 0.1653 0.1650 0.1446 0.0923 0.1761 0.3062 0.3536
CoT 0.2211 0.2218 0.3481 0.2309 0.2143 0.4463 0.3715
LtM 0.1930 0.1918 0.2600 0.2045 0.1752 0.4767 0.3642
MRPP 0.2225 0.2228 0.3641 0.2463 0.2109 0.4350 0.4757
Semantic HaluEval base 0.1075 0.0977 0.1295 0.2224 0.2170 0.1082 0.2587
CoT 0.1507 0.1493 0.1218 0.2018 0.2384 0.1038 0.1468
LtM 0.1828 0.1989 0.1495 0.2231 0.3455 0.1234 0.1492
MRPP 0.1292 0.1329 0.1269 0.2217 0.2551 0.1276 0.1764
Textual PopQA base 0.1507 0.1508 0.1232 0.0992 0.1803 0.2715 0.4548
CoT 0.1342 0.1329 0.1616 0.2019 0.2204 0.1043 0.1204
LtM 0.1718 0.1710 0.2060 0.1233 0.1286 0.2858 0.2076
MRPP 0.1417 0.1391 0.1390 0.2024 0.1834 0.1152 0.1195
Semantic PopQA base 0.1322 0.1274 0.0769 0.1423 0.2190 0.1820 0.4065
CoT 0.1677 0.1657 0.1330 0.0841 0.2075 0.2371 0.1435
LtM 0.1389 0.1363 0.1644 0.1382 0.2567 0.2177 0.1499
MRPP 0.1756 0.1760 0.0848 0.0791 0.2172 0.2429 0.1858
Textual TriviaQA base 0.4652 0.4484 0.4546 0.4893 0.4617 0.2768 0.1262
CoT 0.4976 0.4997 0.3119 0.3898 0.4924 0.3038 0.3421
LtM 0.4427 0.4564 0.3899 0.4210 0.5202 0.2436 0.3003
MRPP 0.4802 0.4802 0.3111 0.3881 0.4868 0.2489 0.3546
Semantic TriviaQA base 0.1657 0.1853 0.0907 0.2557 0.2150 0.1129 0.1556
CoT 0.1598 0.1565 0.1015 0.1674 0.2228 0.1809 0.1515
LtM 0.1060 0.1112 0.0739 0.1364 0.1678 0.1883 0.1406
MRPP 0.1373 0.1384 0.1181 0.1707 0.1564 0.2441 0.1867
Table 20: ECE results of CNN/Daily Mail dataset across different LLMs. The bolded values indicate cases with reduced calibration.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Llama-3.1-8B-Instruct
Textual base 0.1529 0.1487 0.1974 0.1831 0.1882 0.1557 0.2329
CoT 0.1209 0.1227 0.1929 0.2008 0.2377 0.1553 0.1951
LtM 0.1321 0.1198 0.1484 0.1952 0.2605 0.1896 0.1880
MRPP 0.1237 0.1319 0.2409 0.1903 0.2169 0.1658 0.2272
Semantic base 0.1805 0.1753 0.1629 0.1897 0.2506 0.1750 0.1284
CoT 0.1570 0.1568 0.3774 0.1668 0.2370 0.2014 0.2173
LtM 0.1534 0.1510 0.1433 0.1677 0.2636 0.1693 0.2079
MRPP 0.1611 0.1584 0.2984 0.2091 0.2377 0.1626 0.2358
Mistral-7B-Instruct-v0.3
Textual base 0.0995 0.0995 0.1709 0.4466 0.2365 0.1166 0.2240
CoT 0.1536 0.1536 0.2428 0.3138 0.3571 0.1816 0.1310
LtM 0.1850 0.1853 0.2722 0.3012 0.3653 0.1694 0.1188
MRPP 0.1553 0.1553 0.2503 0.3414 0.3042 0.1868 0.1393
Semantic base 0.1391 0.1396 0.1576 0.3481 0.3139 0.1721 0.1273
CoT 0.1646 0.1649 0.3353 0.5001 0.1607 0.1898 0.2619
LtM 0.1757 0.1750 0.4340 0.4852 0.2088 0.1857 0.2578
MRPP 0.1786 0.1780 0.3517 0.5160 0.2510 0.1709 0.2737
Llama-3.1-70B-Instruct
Textual base 0.2342 0.2350 0.2560 0.2136 0.2281 0.1680 0.2037
CoT 0.1495 0.1525 0.2559 0.2533 0.2014 0.2092 0.2457
LtM 0.1484 0.1494 0.2571 0.2168 0.1717 0.1676 0.2456
MRPP 0.1775 0.1754 0.2691 0.2010 0.2011 0.1471 0.2139
Semantic base 0.2467 0.2452 0.2649 0.2374 0.2475 0.1620 0.1010
CoT 0.2011 0.1982 0.2588 0.2665 0.2052 0.1829 0.1847
LtM 0.2005 0.2002 0.2358 0.2731 0.1759 0.1925 0.1757
MRPP 0.2190 0.2138 0.3725 0.2204 0.2160 0.2248 0.1947
Table 21: AUROC results of TruthfulQA. The bolded values indicate cases with reduced performance.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Llama-3.1-8B-Instruct
Info. base 65.37 65.74 62.61 55.57 56.28 61.70 57.74 52.13
CoT 54.87 54.61 54.08 50.22 56.55 53.64 50.82 57.46
LtM 53.69 53.49 66.49 56.37 55.95 57.03 61.21 60.56
MRPP 56.15 55.81 55.76 55.62 63.02 70.01 51.33 80.12
Truth. base 50.35 50.36 54.34 59.17 60.31 54.69 52.43 50.78
CoT 53.06 53.61 51.67 54.79 58.23 53.23 54.84 50.14
LtM 53.16 53.63 53.18 56.15 62.80 52.35 57.10 52.53
MRPP 54.32 54.65 51.22 52.42 55.40 57.67 53.73 50.72
Mistral-7B-Instruct-v0.3
Info. base 51.94 51.91 75.43 63.71 66.54 69.46 62.21 70.23
CoT 76.82 76.82 75.77 62.05 70.74 76.79 52.47 55.91
LtM 74.63 74.63 78.59 69.66 68.55 79.66 59.51 58.12
MRPP 74.23 74.22 69.51 67.95 66.66 80.82 56.80 62.11
Truth. base 54.27 54.26 57.79 56.59 62.20 58.73 55.92 56.20
CoT 53.07 53.06 59.79 59.85 57.53 62.39 51.40 54.98
LtM 55.42 55.41 63.16 59.98 56.10 56.83 50.57 53.26
MRPP 50.43 50.43 57.25 59.71 56.19 55.64 52.88 51.13
Table 22: To mitigate potential biases introduced by the LLaMA-based judge model, we also fine-tuned the evaluation model using Qwen. The evaluation results based on the Qwen judge model are consistent with our previous findings, confirming the robustness of our conclusions.
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Llama-3.1-8B-Instruct
Info. base 56.02 56.52 61.56 54.35 65.94 67.57 72.12 71.36
CoT 54.04 55.31 72.77 53.61 61.02 66.32 51.55 74.78
LtM 59.23 58.75 65.92 50.21 57.74 59.63 64.98 58.63
MRPP 58.6 62.07 67.38 82.95 55.03 52.00 70.88 69.74
Truth. base 56.31 57.33 67.49 56.06 53.19 52.64 75.18 77.3
CoT 56.76 55.54 64.28 52.61 50.32 52.31 60.81 63.94
LtM 58.73 58.66 59.91 53.21 54.24 52.81 56.31 59.4
MRPP 57.75 57.11 66.35 52.55 56.59 52.07 64.59 68.54
Mistral-7B-Instruct-v0.3
Info. base 56.03 56.07 53.91 69.34 78.91 74.14 89.07 85.9
CoT 69.61 69.67 60.44 80.17 74.66 67.95 82.08 53.69
LtM 81.03 80.99 80.95 69.88 63.72 61.17 52.1 64.38
MRPP 75.29 75.29 68.44 81.81 65.99 51.71 74.69 77.45
Truth. base 58.76 58.73 52.31 54.12 62.17 62.23 75.86 75.73
CoT 59.41 59.42 51.59 59.68 64.38 58.91 67.59 65.51
LtM 58.47 58.47 50.22 54.89 59.5 66.14 66.21 68.41
MRPP 58.58 58.58 52.12 59.02 62.68 61.36 64.21 71.65